1. Which of the following is classified as a financing activity on a statement of cash flows?

a. Purchase of a vendor's common stock b. Sale of equipment at a gain c. Payment of dividends to shareholders d. Redemption of a sinking fund

2. Axelton Enterprises sells annual memberships to its shooting lodge. The memberships cost $899 each. On January 1, Axelton sold 8,200 memberships and received cash. What journal entry should Axelton Enterprises make on January 31st if adjusting entries are completed monthly.

a. Debit Unearned Revenue; Credit Cash b. Debit Sales Revenue; Credit Unearned Revenue c. Debit Unearned Revenue; Credit Sales Revenue d. Debit Cash; Credit Sales Revenue

3. Lee Construction enters into a long-term fixed price contract to build an office building for $5,000,000. In the first year of the contract Lee incurs $1,300,000 of cost and the engineers determined that the remaining costs to complete are $2,500,000. Lee billed $2,000,000 and collected $700,000 in year 1. Refer to Lee Construction. How much gross profit should Lee recognize in Year 1 assuming the use of the percentage of completion method? (Round your final answer to the nearest whole dollar.)

a. $410,526 b. $700,000 c. $487,500 d. $1,300,000

4. Gleason Construction enters into a long-term fixed price contract to build an office building for $30,000,000. In the first year of the contract Gleason incurs $8,000,000 of cost and the engineers determined that the remaining costs to complete are $23,000,000 (and indeed they actually were $23,000,000 in the second year). How much gross profit or loss should Gleason recognize in Year 2 assuming the use of the percentage-of-completion method?

a. $1,000,000 profit b. $1,000,000 loss c. $0 d. $741,935 loss
Answer #1
Please give positive ratings so I can keep answering. Thanks!
1. Which of the following is classified as a financing activity on a statement of cash flows?
Answer- c. Payment of dividends to shareholders
a. Purchase of a vendor's common stock It is an investing activity.
b. Sale of equipment at a gain It is an investing activity.
d. Redemption of a sinking fund It is an investing activity.
Answer 2
Annual membership cost              899.00
Number of membership sold          8,200.00
Value 7,371,800.00
To be recognize in January      614,316.67
Journal Entry
Account Debit $ Credit $
Unearned Revenue      614,316.67
Membership fee Income 614,316.67
Answer 3
Lee Construction
Calculation of Gross Profit
Particulars Year 1
Contract Price        5,000,000 A
Cost Incurred to date        1,300,000 B
Cost Incurred during the year        1,300,000 C
Estimated cost to complete        2,500,000 D
Estimated total cost        3,800,000 E=C+B
(Cost Incurred to date+ Estimated cost to complete )
Percentage complete 34.21% F=B/E
(Cost Incurred to date/Estimated total cost)
Revenue to be recognized        1,710,526 G= F*A
Gross Profit/(Loss)            410,526 H=G-C
Option A
Answer 4
Gleason Construction
Calculation of Gross Profit
Particulars Year 1 Year 2
Contract Price      30,000,000 30,000,000 A
Cost Incurred during the year        8,000,000 23,000,000 C
Cost Incurred to date        8,000,000 31,000,000 B
(Cost Incurred of the present year plus Cost incurred of the previous year)
Estimated cost to complete      23,000,000 D
Estimated total cost      31,000,000 31,000,000 E=B+D
(Cost Incurred to date+ Estimated cost to complete )
Percentage complete 25.81% 100.00% F=B/E
(Cost Incurred to date/Estimated total cost)
Revenue to be recognized        7,741,935 22,258,065 G
(Contract Price*Percentage complete)- revenue previously recognized
Year 1: 25.81% completed. Revenue recognized = 25.81% x $ 30,000,000 = $ 7,741,935.
Year 2: 100% completed. Revenue recognized = 100% x $ 30,000,000 – $ 7,741,935 (previously recognized) = $ 22,258,065.
Gross Profit/(Loss)          (258,065)       (741,935) H=G-C
Option D